Two spheres A and B are placed in the arrangement shown below.
(a) If mA = 5m and mB = 7m, where on the dashed line should a third sphere C of mass 7m be placed so that the net force on it is zero?
at the midpoint between A and B, to the right of A , to the left of B, between A and B, closer to B, between A and B, closer to A

(b) If the distance between the two spheres A and B is 500 cm, find the location for the third sphere C so that the net force on it is zero.
cm to the right of sphere B
Answer #1

A)

Force, F = GMm/r^2

G ma mc/rca2 = G mb mc/rcb2

rbc/rca = sqrt(mb/ma) = sqrt(7/5) = 1.18

rbc = 1.18 rac

Sphere C must be placed between the spheres B and A and closer to sphere A

B)

rbc/(500 - rbc) = sqrt(7/5) = 1.18

rbc = 591.61 - 1.18 rbc

2.18 rbc = 591.61

rbc = 271.38 cm

Location of sphere C is 271.38 cm from sphere B

Comment in case any doubt please rate my answer....